OpenStudy (anonymous):

To answer the riddle, write the numbers in order from least to greatest. Then write the corresponding letters in the same order. HELP PLEASE!! GIVING METALS! -1 1/4 : A -2.05 : E -1/5 : T 3/4 : N 0 : I -2.7 : S -5/6 : R -1 1/8 : S 0.95 : L -0.6 : A -1.3 : W 0.8 : A -2 1/2 : H 1/8 : O Riddle : _______________

jhonyy9 (jhonyy9):

-1 1/4 = ? can you rewrite it like a fraction ?

OpenStudy (anonymous):

idk :c

jhonyy9 (jhonyy9):

-5/4 do you know how i have got it

jhonyy9 (jhonyy9):

- 1 1/4 = - (1*4+1)/4 = - (4+1)/4 = - 5/4 ok ?

OpenStudy (anonymous):

i just want them in order so i can get this over with :c plz

jhonyy9 (jhonyy9):

so for you can getting this answer right you need to transform every fractions with integer like this and after just divide it for get the number result like 1,002 for example ok ?

OpenStudy (anonymous):

i dont get it

jhonyy9 (jhonyy9):

so than A. will be how many ?

OpenStudy (anonymous):

-5/4? but how do i put it in least to greatest so i can get the riddle

jhonyy9 (jhonyy9):

A. will be -5/4 so this mean 5 divide by 4 5/4 = 1,25 4 -- 10 8 --- 20 20 --- 00 so than A. will be - 1,25 than continue with B.

OpenStudy (anonymous):

I would suggest you find the smallest number in that group.

jhonyy9 (jhonyy9):

E. is -2,05 so than E is greater or smaler than A ?

OpenStudy (anonymous):

omg :( i just neeed the riddle!

OpenStudy (anonymous):

pleaseeeeee

OpenStudy (anonymous):

Sounds like you just want someone to give you the answer...that is not the point of this website...perhaps you should go to another website.

jhonyy9 (jhonyy9):

exactly

OpenStudy (anonymous):

pathetic.

OpenStudy (anonymous):

Yes the fact that you do not want to learn how to do this is pathetic.

OpenStudy (anonymous):

i dont get the way you're teaching me.

OpenStudy (anonymous):

I think that is exactly what we are trying to do..Let's try this...Which of those numbers is the smallest...just find the smallest number.

OpenStudy (anonymous):

-0.8 ??

OpenStudy (anonymous):

No... remember when you are looking at negative numbers the "larger" the number looks the smaller it actually is. so in other words. -5 is smaller than -1.

OpenStudy (anonymous):

So with that in mind what looks to be the smallest number in that group?

OpenStudy (anonymous):

mmm, -2.7 ???

OpenStudy (anonymous):

Yes! So now you know which letter is the start of the riddle! Now what is the next smallest?

OpenStudy (anonymous):

-2.05 ??

OpenStudy (anonymous):

Well I see two possibilities -2.05 and -2 1/2, how do you write -2 1/2 as a decimal?

OpenStudy (ybarrap):

You can cut this an paste into excel, lining up numbers and letters. Sort by numbers and your answer will be magically revealed.

OpenStudy (anonymous):

-2 1/2 = -2.5 ? so -2 1/2 is next. right?

OpenStudy (anonymous):

You got it. Now what's the next smallest? (You're cooking now)

OpenStudy (anonymous):

-2.05 ?

OpenStudy (anonymous):

Good let's keep going...

OpenStudy (anonymous):

-1.3 ?

OpenStudy (anonymous):

Good ... keep going

OpenStudy (anonymous):

-1 1/4 = -1.25 ? that's next right?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Yep great job... keep it up... let 's look at what the letter order is so far before we go any further.

OpenStudy (anonymous):

s h e w a , i got that so far

OpenStudy (anonymous):

It is looking good so far since the letter order looks like it might be some kind of answer...maybe " She was..." ... Just curious what was the riddle that this is supposed to answer?

OpenStudy (anonymous):

i dont know. my teacher gave it to us for homework

OpenStudy (anonymous):

Oh because it looks like we are coming up with the answer to a riddle.

OpenStudy (anonymous):

Not the riddle itself

OpenStudy (anonymous):

Oh, "Why did Francis Fraction want to become a psychologist"

OpenStudy (anonymous):

I think we are on the right track then. Keep going.

OpenStudy (anonymous):

next is -1 1/8 = 1.125

OpenStudy (anonymous):

Yep and that is a good sign since we now actually have "She was"! Don't stop now.

OpenStudy (anonymous):

-5/6 = -0.833.. ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

-0.6 ?

OpenStudy (anonymous):

Yes, I think you understand this now. Way to go!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem. Let me know if you need any more help...and oh yeah ... watch your language, young lady... ;-)